OpenStudy (anonymous):

There are 7 girls in the bus. Each girl has 7 backpacks. Each backpack has 7 big cats. Every big cat have 7 little cats. How many legs are there in the bus?

OpenStudy (anonymous):

4*7=28 for each backpack 28*7=196 for each girl 196*7=1372 Total BIG cat legs Next, we'd have to find the number of girl's legs. 7*2=14. Add it up: 9604+1372+14=10,990 There you are. 10,990 is the total number of legs. Read more: